Question 18

Calculate the missing value for the probability P(Z <?) = 0.5
(1) 1
(2) -1
(3) 0
(4) 0.5
(5) -0.5
1
Expert's answer
2020-03-02T14:52:17-0500

"P(Z<0)=\frac{1}{\sqrt{2\pi}}\int\limits_{-\infty}^0e^{-\frac{x^2}{2}}dx=\frac{1}{2}"

Answer: (3) .